OpenStudy (anonymous):

Integral x/ (1+esinx) ....from 0 to pie/2

OpenStudy (ash2326):

Is this your question \[\int_{0}^{\pi/2} \frac{x}{1+e^x \sin x} dx \] ?

OpenStudy (anonymous):

Yes...

OpenStudy (turingtest):

wolf won't do the indefininite, and gives some random-looking decimal for the answer to the definite

OpenStudy (anonymous):

Oh wait...I just checked the question again...its just "e" . Not e^x .

OpenStudy (turingtest):

doesn't seem to help much wolf still refuses to help

OpenStudy (turingtest):

this would lead me to think that this cannot be represented in terms of elementary functions http://www.wolframalpha.com/input/?i=integral+x%2F%281%2Be*sinx%29dx though perhaps for the definite we can shift the bounds somehow, and use some even-odd properties to handle it I have to get to class, good luck!

OpenStudy (anonymous):

Ok...thanks! :)